DEF Company has provided the following information: Sales price per unit $52 Variable cost per unit...

  • DEF Company has provided the following information:

Sales price per unit

$52

Variable cost per unit

16

Fixed costs per month

$14,000

Calculate the contribution margin per unit.
Answer #1

Contribution Margin (%) = (Selling Price - Variable Cost) / Selling price * 100

Ie (52-16)/52*100 = 69.23%

In amount Contribution = 52-16 = 36.